Show your solutions:


The speed of light in a certain material is 2 x 108 m/s. The speed of light in a vacuum is 3 x 108 m/s. Calculate the index of refraction of the material.


1
Expert's answer
2021-06-01T10:49:19-0400

We can find the index of refraction of the material as follows:


"v=\dfrac{c}{n},""n=\dfrac{c}{v}=\dfrac{3\cdot10^8\ \dfrac{m}{s}}{2\cdot10^8\ \dfrac{m}{s}}=1.5"